# Long Context Scaling: Divide and Conquer via Multi-Agent Question-driven Collaboration

### v1

@@ 4.4 Comparison with Other Multi-Agent Frameworks
Table 3: Comparison with other agent-based works on GraphWalks and MRCR. XpandA significantly outperforms LongAgent and CoA in both success and progress rates on GraphWalks and MRCR, demonstrating superior efficiency in information acquisition and structured reasoning for multi-hop long-context problems.
@@
success progress/steps success progress/steps success progress/steps
@@ 5.1 Diverse Chunking Strategies Meet Context of Vastly Varying Length
Figure 3: Impact of different chunking strategies on Xpanda in dataset MRCR (2 Needle). Optimal static performance occurs when the unit chunk size is $[\frac{1}{16},\frac{1}{4}]$ of the input length, while dynamic chunking controls the proportion within this range and achieves 18% average improvement.
We analyze the impact of chunking strategies on XpandA’s performance, shown in Figure 3, via ablation experiments on the MRCR (2 Needles) benchmark (16k–1M sequences), evaluating retrieval success rate. Static chunking (2k–128k) shows that when the chunk size exceeds the input length, XpandA degenerates to a single-agent system, suffering full-context-like performance decay. Conversely, overly small chunks introduce excessive agents, collecting irrelevant noise and degrading accuracy. Optimal static performance occurs when unit chunk size is $[\frac{1}{16},\frac{1}{4}]$ of the input sequence length.
@@ 5.2 LLM Reasoning Scheme Ablation
To further explore the effectiveness of the question-driven scheme in xpanda, we systematically remove and modify the component from XpandA to observe the performance changes. Plan-and-Solve [78] and Fact-and-Reflection [103] can help improve the LLM reasoning performance through the management of atomic plans or atomic facts, behaving just like the subquestion state tracking in XpandA, and thus can replace the planning scheme in XpandA without reconstruction of pipeline. We implement Directly Answer by setting the question tracker $P$,$T$ to null value and gathering all Explorers’ direct answers to be chosen by Decider. As shown in Figure 4(a), XpandA outperforms other reasoning schemes by 4–12% in F1 score of NarrativeQA, which shows that the inclusion of question-driven scheme in XpandA efficiently reinforces the multi-agent workflow by providing instruction that is more explicit to identify and explore the context and mitigates the inter-agent misalignment.
@@ 5.4 Empirical Inference Time Analysis
Figure 5: XpandA scheme exhibits 1.5x speedup in inference time compared to SOTA long context language model on long texts.
We analyze the inference time of XpandA on 4×A100_80G GPUs (Figure 5), using Qwen2.5:7B-Instruct-1M as the backbone to avoid API-related latency. For comparison, we include Llama3.1-Instruct-1048k [62] in full-context mode. XpandA with dynamic chunking outperforms all baselines, achieving a 1.5× speedup over Qwen2.5-1M (with SOTA sparse attention optimization [91]) while maintaining linear complexity with fixed chunk sizes. In contrast, traditional full-context models (limited to 256k context on 4×A100) are significantly slower. Dynamic chunking also balances inference time and model-call latency better than static strategies (2k/8k) used in CoA [100] and LongAgent [101], which suffer from accumulated per-agent latency—especially critical for commercial API calls.
@@ 6 Conclusion
7 Limitations
Although XpandA provides an excellent paradigm for addressing the weak constraints and operational inefficiencies of Long Context Multi-Agent Systems in prior research, future directions can address the following limitations to further enhance its capabilities and efficiency. First, the scalability of XpandA on models with weak instruction-following abilities could be improved through in-context learning, simplifying workflow instructions, or exploring more efficient workflows. Second, the current experiments limit input sequence lengths to under 1M; future work could further investigate XpandA’s behavior and performance on longer sequences.
@@ Appendix A Proof of Completeness
Hypothetical Premise: the LLM($I$,$\theta$) acting an agent can complete the problem decomposition and reasoning of a specified context chunk within k tokens context window.
@@
The Explorer accepts $m_{x,y}$=1 and then 2, and so on until all chunks are received. Specifically, in ($y-1$) rounds of scan direction reverse, the Explorer would accept all items in M.
@@
Scan from left to right. For every y chunks, at least one chunk is accepted. So all items in M can be accepted in $y$ scan, namely $y-1$ replay. It’s not non-trivial, for:
$\forall x,DrA\ accept\ M.$ (19)
Appendix B Supplementary Experiments
(a) Full model
(b) RAG model
(c) XpandA
Figure 6: Dual-Needle-in-a-Haystack test. (a) Full model. (b) RAG model. (c) XpandA.
The Dual-Needle-in-a-Haystack test is designed to rigorously evaluate a model’s capability to concurrently identify, associate, and utilize two distinct pieces of information ("needles") embedded within a voluminous and distractor-rich textual corpus ("haystack"). This task extends beyond single-needle retrieval by imposing a more significant challenge on the model’s long-range dependency understanding, information integration, and sustained attention across intensive contexts. This appendix presents a detailed performance analysis of the Full-Context, RAG, and XpandA methodologies on this task.
The "Length of Haystack", represented on the X-axis of Figure 6, was varied across discrete token counts from 1k to 128k. Additionally, two distinct "needles" were synthetically generated for each trial, whose depth distribution is represented on the Y-axis of the heatmaps. For a given trial, the two needles were semantically independent but both were required to answer the query. The "Depth Percent of Needles", shown on the Y-axis of the heatmaps, indicates the predefined contextual zones where the needles were inserted.
Figure 6(a) presents the performance heatmap for the Full-Context method. The overall average accuracy achieved was 52.58%. The general trend of degrading accuracy is observed as the haystack length increases beyond approximately 64k tokens, indicating the increasing difficulty in processing, retaining, and associating information from both needles when the context window is saturated. Successfully retrieving and associating two needles becomes more challenging when one or both are deeply embedded and when their relative distance is large within an extensive context.
The RAG method in Figure 6(b) achieved an overall average accuracy of 64.01%. As the heatmap indicates, performance is sensitive to haystack length. While RAG’s retrieval mechanism improves overall accuracy compared to the Full-Context method (Figure 6(a)), it exhibits limitations in effectively locating and leveraging information in increasingly longer documents. Its effectiveness negatively correlates with document length, suggesting potential retrieval challenges in complex long-context scenarios.
Our proposed multi-agent framework, XpandA in Figure 6(c), achieved an overall average accuracy of 72.53%, representing a significant improvement over the RAG (64.01%) and Full-Context (52.58%) methods on this task. The heatmap provides compelling visual evidence of XpandA’s enhanced robustness, effectively mitigating the substantial performance degradation typically observed in baseline approaches, particularly under challenging conditions such as extended context lengths, deep needle locations, and considerable relative distances between needles.

### v3

@@ 4.4 Comparison with Other Multi-Agent Frameworks
Table 3: Comparison with other agent-based works on GraphWalks and MRCR. XpandA significantly outperforms LongAgent and CoA in both success (suc), progress rates (prog) and agent invoke steps (steps) on GraphWalks and MRCR, demonstrating superior efficiency in information acquisition and structured reasoning for multi-hop long-context problems.
@@
suc prog/steps suc prog/steps suc prog/steps
@@ 5.1 Diverse Chunking Strategies Meet Context of Vastly Varying Length
We analyze the impact of chunking strategies on XpandA’s performance, shown in Figure 3, via ablation experiments on the MRCR (2 Needles) benchmark (16k–1M sequences), evaluating retrieval success rate. Static chunking (2k–128k) shows that when the chunk size exceeds the input length, XpandA degenerates to a single-agent system, suffering full-context-like performance decay. Conversely, overly small chunks introduce excessive agents, collecting irrelevant noise and degrading accuracy. Optimal static performance occurs when unit chunk size is $\frac{1}{16},\frac{1}{4}]$ of the input sequence length.
@@
Figure 3: Impact of different chunking strategies on Xpanda in dataset MRCR (2 Needle). Optimal static performance occurs when the unit chunk size is $\frac{1}{16},\frac{1}{4}]$ of the input length, while dynamic chunking controls the proportion within this range and achieves 18% average improvement.
@@ 5.2 LLM Reasoning Scheme Ablation
To evaluate XpandA’s question-driven scheme, we modified components and observed performance, shown in Figure 4(a). Plan-and-Solve (Wang et al., 2023a) and Fact-and-Reflection (Zhao et al., 2024c), similar to XpandA’s subquestion state tracking, can replace its planning scheme. We also implemented Directly Answer by nullifying question trackers $P,T$ and having the Decider choose from Explorers’ direct answers. Figure 4(a) shows XpandA outperforms these schemes by 4–12% in F1 score on NarrativeQA. This indicates XpandA’s question-driven scheme effectively enhances the multi-agent workflow by providing explicit instructions for context identification and exploration, mitigating inter-agent misalignment.
@@ 5.4 Empirical Inference Time Analysis
Figure 5: XpandA scheme exhibits 1.5x speedup in inference time compared to SOTA long context language model on long texts. The x-axis represents the length of the context, and the y-axis represents the inference time (seconds).
We analyze the inference time of XpandA on 4×A100_80G GPUs (Figure 5), using Qwen2.5:7B-Instruct-1M as the backbone to avoid API-related latency. For comparison, we include Llama3.1-Instruct-1048k (Pekelis et al., 2024) in full-context mode. XpandA with dynamic chunking outperforms all baselines, achieving a 1.5× speedup over Qwen2.5-1M (with SOTA sparse attention optimization (Yang et al., 2025)) while maintaining linear complexity with fixed chunk sizes. In contrast, traditional full-context models (limited to 256k context on 4×A100) are significantly slower. Dynamic chunking also balances inference time and model-call latency better than static strategies (2k/8k) used in CoA (Zhang et al., 2024d) and LongAgent (Zhao et al., 2024a), which suffer from accumulated per-agent latency—especially critical for commercial API calls.
@@ 6 Conclusion
7 Impact Statement
This paper presents a framework for processing long context based on Multi-Agent System, without involving any social or ethical issues or consequences.
8 Acknowledgements
This work is supported by funding from the National Natural Science Foundation of China (NSFC No.62161160339).
@@ Appendix A Proof of Completeness
[…]
@@
The Explorer accepts $m_{x,y}$=1 and then 2, and so on until all chunks are received. Specifically, in (y1y-1) rounds of scan direction reverse, the Explorer would accept all items in M.
@@
Scan from left to right. For every y chunks, at least one chunk is accepted. So all items in M can be accepted in $y$ scan, namely y1y-1 replay. It’s not non-trivial, for:
$\forall x,XpandA\ accept\ M.$ (19)